OpenStudy (anonymous):

solve 3(y – 5) + 2 = 5 step by step!!

OpenStudy (anonymous):

multiply 3(y-5) we get 3y - 15 +2=5. add -15+2=-13. add-13 to both sides. we get 3y=18 divide 3 on both sides we get y=6

OpenStudy (anonymous):

3y-3x5+2=5 3y-15=5-2 3y=15+3 3y=18 y=18/3 y=6